OpenStudy (anonymous):

evaluate the indefinite integral and explain how to do it please!! (3t+2)^(2.4)dt

OpenStudy (dumbcow):

Use u substitution u = 3t+2 du = 3dt \[=\frac{1}{3} \int\limits_{}^{}u^{2.4} du\]